OpenStudy (landon):

Find the radius of a circle in which an inscribed square has a side of 8 inches. x =

OpenStudy (eliesaab):

Use this to find D=2r , knowing that s=8 \[ D^2= 2 s^2 \] Then find \[ r=\frac D2\]

OpenStudy (michaelbp):

4sqrt2

OpenStudy (mrnood):

 D is the diagonal of the square You can work it out from pythagoras - or use the well known fact that diagaonal of square = side * sqrt(2) (This is the same as @eliesaab said above - in a slightly different way) You can also see that D is the diameter of the circle - so the radius is D/2